# Diversity Enhances an LLM’s Performance in RAG and Long-context Task

### v1

@@ Abstract
Figure 1: For both Q&A and summarization tasks, the initial dataset is divided into sentences or chunks, and corresponding embeddings are extracted. In a traditional pipeline, query embeddings are generated and used to select relevant materials to LLMs for downstream tasks. In contrast, methods like MMR and FPS incorporate diversity in a greedy manner when selecting relevant sentences. This approach increases the likelihood of including the correct answer within the chosen sentences or chunks.
@@ 1 Introduction
Various strategies have been devised to address the limited context window issue in LLMs. Longformer [7] applies attention to immediate local neighbors, reducing the time complexity from $O(N^{2})$ to $O(NM)$, with $M$ representing the considered neighbors. This approach, however, necessitates significant alterations to the attention mechanism, which is not commonly adopted in contemporary LLMs such as GPT[3], Llama[8], and Gemini[9]. An alternative strategy is to expand the context window at inference [10]. Although this can mitigate the modification during the training process, it still demands changes to the attention architecture during the inference time, which is not accessible for close-source models like GPT.
Several strategies have been proposed to address the limited context window in LLM from the training perspective. The Longformer model [7] employs attention mechanisms focused on immediate local neighbors, reducing the time complexity from $O(N^{2})$ to $O(NM)$, where $M$ denotes the number of neighbors considered. However, this method requires substantial modifications to the attention mechanism, which are not widely adopted by contemporary LLMs such as GPT [3], LLaMA [8], and Gemini [9]. Another approach involves extending the context window during inference [10]. While this mitigates the need for training-time modifications, it necessitates changes to the attention mechanism at inference and full access to the model architecture—an obstacle for closed-source models like GPT.
Previous methods primarily focus on modifying LLMs to increase their context window. However, a more straightforward approach is to first select the most relevant documents while ensuring they fit within the LLM’s context window. For a given query, multiple documents are split into smaller chunks or sentences. The embeddings for both the query and the split documents are then computed. Similarity metrics, such as cosine similarity or Euclidean distance, are subsequently used to identify the most relevant sentences.
However, relying solely on the similarity between a query and segmented documents can result in overlooking critical information due to excessive focus on similar content. Previous studies have introduced greedy algorithms, such as MMR [11] and FPS [12], to improve diversity during the selection process. Related work introduced Hypothetical Document Embedding (HyDE) and LLM reranking to enhance diversity in Q&A tasks, claiming their method outperforms MMR [13, 14]. However, these studies did not address the recall of relevant documents prior to LLM generation, which is more pertinent to diversity considerations. Additionally, they did not explore various hyperparameters within MMR. Then, they have not explored the impact of reordering of selected sentences or chunks on the downstream tasks. In this paper, we aim to address this gap by conducting experiments to demonstrate the significance of diversity in long context summarization and RAG-based Q&A tasks at multiple levels: sentence-level for single documents, chunk-level across entire datasets, and sentence-level in summarization.
@@ 2 Methodology
2.1 MMR and FPS for Diversity
@@
The concept of MMR involves selecting a subset $S$ from a large dataset $T$ [11]. MMR uses a greedy algorithm that starts with the selected set $S$ being empty and the remaining set $R$ being the entire dataset $T$. In each iteration, an element is chosen based on a locally optimal selection process, as defined in Eq. 1. The parameter $\alpha$ balances the trade-off between rewards and diversity. Let $r_{i}$ denote the reward of the $i$-th item, and $cos(i,j)$ represent the cosine similarity between the $i$-th and $j$-th items in the selected subset. $W$ is a subset of $S$ that includes the most recently selected examples, reducing the emphasis on earlier selections. For example, if $w=10$, $W$ consists of the last 10 selected samples from $S$, while all previously selected examples are excluded from diversity considerations. The objective of MMR is to maximize rewards while ensuring sufficient diversity among the selected items. This iterative process continues until a termination criterion, such as reaching a predefined maximum number of tokens, is met.
@@
FPS
The concept of FPS originates from the field of 3D computer vision [12]. Its primary goal lies in selecting a diverse set of points from a given point cloud, which aids in hierarchical feature extraction for downstream applications. The process begins with a randomly selected initial point. In each subsequent iteration, a new point is chosen based on its distance from all previously selected points. When comparing FPS to MMR, we find that both are greedy methods that promote diversity by selecting points that differ from those chosen. However, FPS does not incorporate the concepts of a context window or reward. If we modify FPS to include these elements, the modified FPS will be equivalent to MMR, with the key difference being that MMR uses cosine similarity, while FPS relies on Euclidean distance for measuring similarity.
2.2 Combine MMR and FPS with LLM for Diversity on Q&A and Summarization
Extending MMR and FPS techniques for LLMs in tasks such as Q&A and summarization is relatively straightforward as shown in Fig. 1. These techniques employ a greedy approach to iteratively balance the similarity of selected sentences or chunks to the query with the diversity among the selected sentences or chunks. This method enhances the likelihood of selecting the most relevant sentences or chunks for LLMs in downstream tasks. Lastly, inspired by [15], a heuristic rearrangement scheme is implemented to enhance the likelihood of identifying the correct answer from the retrieved documents.
@@ 2.2 Combine MMR and FPS with LLM for Diversity on Q&A and Summarization
To evaluate the ability of LLMs on accurately extracting the correct answer, a query, a document, and a corresponding answer are initially provided. Documents are pre-processed by dividing them into sentences or chunks, and their embeddings are extracted beforehand. Both the query and the segmented documents are processed using encoder-only models to generate embeddings. In MMR, similarity is measured using the cosine angle, whereas in FPS, Euclidean distance is used to assess similarity. For benchmarking Q&A performance, two metrics should be evaluated:
@@
Figure 2: The impact of different hyperparameters: $\alpha$, $w$, $c_{r}$ on the recall of the Natural Question dataset of single document Q&A. The first and second subfigures illustrate the recall ratios of answers contained in the selected documents for SB+MMR and SB+FPS. When the weight parameter $w=1$, they are equivalent to SB. From the results, we can conclude that both SB+MMR and SB+FPS outperform SB. The last two subfigures display the latency of SB+MMR and SB+FPS. SB+FPS shows slightly worse performances than SB+MMR, and the latency of SB+MMR is significantly lower, especially when the context window is very long. Considering these two aspects, SB+MMR is more suitable for practical use compared to SB+FPS.
@@ 3.1 Single Document Q&A
For single document Q&A, three datasets are included: 1. Natural Question [18], 2. Trival QA [19] and 3. Narrative QA [20]. For each dataset, it is composed of thousands of (query, document, answer) pairs where the answer exists within the document and answers the query. For each document, we split it into sentence using Spacy package [21]. SentenceBERT (SB) is utilized as the encoder to extract embeddings from sentences in different experiments [22]. Then, different compression ratios, i.e., $c_{r}=0.05,0.1,0.2$ are utilized. Here, $c_{r}=0.05$ represents that the fraction of the number of selected tokens over the total number of tokens should be 0.05, i.e., the termination condition when 0.05 of all tokens are selected for answering the query. As for $\alpha$ and $w$, different hyperparameters are tested in a two-level iteration. The first coarse-level iteration utilizes $\alpha$ from [0, 0.25, 0.5, 0.7, 0.8, 0.9, 0.95, 1] and $w$ from [0, 10, 100, 1000, 1000000]. Then, the best performing from the first coarse-level iteration is selected. For the second granular-level iteration, it further divides the neighbors of the best performing first coarse-level hyperparameters and select the ones that have the best performance. An example of the best performing hyperparameters in the Natural Question dataset in the coarse level can be found in Fig. 2. In particular, for Natural question, $\alpha=0.55$ and $w=1$ is the best for MMR and $\alpha=0.5$ and $w=1$ is the best for FPS. For Narrative Q&A, $\alpha=0.55$ and $w=5$ is the best for both MMR and FPS. For Trival Q&A, $\alpha=0.6$ and $w=3$ is the best for MMR and $\alpha=0.7$ and $w=1$ is the best for FPS.
@@
Based on the results across various datasets, we can assert that diversity significantly enhances the recall of the correct answer within the selected document, as demonstrated in Table 1, showing an improvement of 2% to 5%. When the extracted sentences are summarized by GPT4 using the prompt shown in Figure 3, the advantages of SB+MMR and SB+FPS over SB alone remain evident, as shown in Table 2. Additionally, we observe that the performance of Trivial Q&A after LLM is better than the retrieved sentences, with a consistent result of approximately 76%. This suggests that the performance is largely influenced by the LLM, possibly due to pretraining on Trivial Q&A, even when the retrieved documents are provided. FPS, using distance as the evaluation metric, performs slightly worse than MMR, which uses cosine similarity. Moreover, MMR is faster than FPS because computing cosine similarity is quicker than Euclidean distance in Python, especially as the compression ratio increases, as shown in Figure 2. This conclusion generally holds true across different datasets. The speed advantage of MMR becomes more critical as the number of candidates increases with the dataset size. Consequently, MMR will be used in the multiple document comparison in the next section.
@@ 5 Limitation
There are several limitation on this works. To begin with, we only work on English dataset, while multilingual datasets should be tested to prove the importance of diversity on other language. In addition, this work focuses on research dataset while more work is supposed to be conducted on industrial datasets. Lastly, for extremely large dataset, more engineering work on parallelization like tree structures should be conducted to reduce latency.
Appendix A Example Appendix
A.1 Prompts for Q&A and Summarization

### v2

@@ Abstract
Figure 1: For both Q&A and summarization tasks, the initial dataset is divided into sentences or chunks, and corresponding embeddings are extracted. In a traditional pipeline, query embeddings are generated and used to select relevant materials to LLMs for downstream tasks. In contrast, methods like MMR and FPS incorporate diversity in a greedy manner when selecting relevant sentences. This approach increases the recall of the correct answer within the chosen sentences or chunks.
@@ 1 Introduction
Various strategies have been devised to address the limited context window issue in LLMs. Longformer [7] applies attention to immediate local neighbors, reducing the time complexity from $O(N^{2})$ to $O(NM)$, with $M$ representing the considered neighbors. An alternative strategy is to expand the context window at inference [8]. Lastly, a more straightforward approach is to first select the most relevant documents while ensuring they fit within the LLM’s context window. For a given query, multiple documents are split into smaller chunks or sentences. The embeddings for both the query and the split documents are then computed. Similarity metrics, such as cosine similarity or Euclidean distance, are subsequently used to identify the most relevant sentences.
However, relying solely on the similarity between a query and segmented documents can result in overlooking critical information due to excessive focus on similar content. Previous studies have introduced greedy algorithms, such as MMR [9] and FPS [10], to improve diversity during the selection process. In this paper, we aim to demonstrate the significance of diversity in long context summarization and RAG-based Q&A tasks at multiple levels: sentence-level for single documents, chunk-level across entire datasets, and sentence-level in summarization.
@@ 2 Methodology
2.1 FPS and MMR for Diversity
FPS
The concept of Farthest Point Sampling (FPS) originates from the field of 3D computer vision [10]. Its primary goal is to select a diverse set of points from a given point cloud, which facilitates hierarchical feature extraction for downstream applications. The FPS process begins with a randomly selected initial point. In each subsequent iteration, a new point is chosen based on its maximum distance from all previously selected points as shown in Eq. 1 where $S$ refers to the selected set, $R$ refers to the remaining set and $dist(i,j)$ represents the Euclidean distance between $i$ and $j$. For initialization, the selected set $S$ is empty and the remaining set $R$ equals to the entire dataset. An illustration of FPS is provided in Figure 2, where 50 points are selected from a set of 1000 points in a 2D domain. Unlike some other selection methods, FPS solely relies on Euclidean distance to promote diversity among the selected points.
Figure 2: An example of FPS: selecting 50 points from 1000 points in 2D space. red: currently selected point; green: previous selected points; blue: unselected points
Figure 3: The impact of different hyperparameters: $\alpha$, $w$, $c_{r}$ on the recall accuracy and latency using FPS of the Natural Question dataset of single document Q&A.
@@ 2.1 FPS and MMR for Diversity
In contrast, Maximal Marginal Relevance (MMR) selects a subset $S$ from a large dataset with the dual objectives of maximizing rewards and ensuring diversity [9]. At each iteration, an element is chosen based on a locally optimal criterion defined in Eq. 2. Here, the reward $r_{i}$ of the $i$-th item is balanced against the similarity measure, where $\text{cos}(i,j)$ denotes the cosine similarity between the $i$-th item and those in a context window $W$ (comprising the most recently selected examples). This context window, for example, could include the last 10 selected samples when $w=10$, reducing the influence of earlier selections. The iterative process continues until a predefined termination criterion, such as reaching a maximum number of tokens, is met.
@@
When comparing the two methods, both FPS and MMR use greedy selection to promote diversity. The key differences lie in their selection criteria: while FPS focuses exclusively on maximizing Euclidean distance for diversity, MMR introduces an additional reward factor and employs cosine similarity within a dynamic context window to balance between relevance and diversity. In this paper, we extend FPS by incorporating the reward mechanism and the dynamic context window. Ultimately, the key distinction between the modified FPS and MMR lies in their similarity metrics: the former uses Euclidean distance, while the latter relies on cosine similarity.
2.2 Combine MMR and Modified FPS with LLM for Diversity on Q&A and Summarization
Extending MMR and Modified FPS techniques for LLMs in tasks such as Q&A and summarization is relatively straightforward as shown in Fig. 1. These techniques employ a greedy approach to iteratively balance the similarity of selected sentences or chunks to the query with the diversity among the selected sentences or chunks. This method enhances the likelihood of selecting the most relevant sentences or chunks for LLMs in downstream tasks. Lastly, inspired by [11], a heuristic rearrangement scheme is implemented to enhance the likelihood of identifying the correct answer from the retrieved documents.
@@ 2.2 Combine MMR and Modified FPS with LLM for Diversity on Q&A and Summarization
To evaluate the ability of LLMs on accurately extracting the correct answer, a query, a document, and a corresponding answer are initially provided. Documents are pre-processed by dividing them into sentences or chunks, and their embeddings are extracted beforehand. Both the query and the segmented documents are processed using encoder-only models to generate embeddings. In MMR, similarity is measured using the cosine angle, whereas in Modified FPS, Euclidean distance is used to assess similarity. For benchmarking Q&A performance, two metrics should be evaluated:
@@ 3.1 Single Document Q&A
For single document Q&A, three datasets are included: 1. Natural Question [14], 2. Trival QA [15] and 3. Narrative QA [16]. For each dataset, it is composed of thousands of (query, document, answer) pairs where the answer exists within the document and answers the query. For each document, we split it into sentence using Spacy package [17]. SentenceBERT (SB) is utilized as the encoder to extract embeddings from sentences in different experiments [18]. Then, different compression ratios, i.e., $c_{r}=0.05,0.1,0.2$ are utilized. Here, $c_{r}=0.05$ represents that the fraction of the number of selected tokens over the total number of tokens should be 0.05, i.e., the termination condition when 0.05 of all tokens are selected for answering the query. As for $\alpha$ and $w$, different hyperparameters are tested in a two-level iteration. The first coarse-level iteration utilizes $\alpha$ from [0, 0.25, 0.5, 0.7, 0.8, 0.9, 0.95, 1] and $w$ from [0, 10, 100, 1000, 1000000]. Then, the best performing from the first coarse-level iteration is selected. For the second granular-level iteration, it further divides the neighbors of the best performing first coarse-level hyperparameters and select the ones that have the best performance. An example of the best performing hyperparameters for FPS on recall accuracy and latency in the Natural Question dataset in the coarse level can be found in subfigure (a) of Fig. 3. More details on the impact of hyperparameters can be found in Fig. 6. In particular, for Natural question, $\alpha=0.55$ and $w=1$ is the best for MMR and $\alpha=0.5$ and $w=1$ is the best for FPS. For Narrative Q&A, $\alpha=0.55$ and $w=5$ is the best for both MMR and FPS. For Trival Q&A, $\alpha=0.6$ and $w=3$ is the best for MMR and $\alpha=0.7$ and $w=1$ is the best for FPS.
@@
Based on the results across various datasets, we can assert that diversity significantly enhances the recall of the correct answer within the selected document, as demonstrated in Table 1, showing an improvement of 2% to 5%. When the extracted sentences are summarized by GPT4 using the prompt shown in Figure 4, the advantages of SB+MMR and SB+FPS over SB alone remain evident, as shown in Table 2. Additionally, we observe that the performance of Trivial Q&A after LLM is better than the retrieved sentences, with a consistent result of approximately 76%. This suggests that the performance is largely influenced by the LLM, possibly due to pretraining on Trivial Q&A, even when the retrieved documents are provided. FPS, using distance as the evaluation metric, performs slightly worse than MMR, which uses cosine similarity. Moreover, MMR is faster than FPS because computing cosine similarity is quicker than Euclidean distance in Python, especially as the compression ratio increases, as shown in subfigure (b) of Figure 3. This conclusion generally holds true across different datasets. The speed advantage of MMR becomes more critical as the number of candidates increases with the dataset size. Consequently, MMR will be used in the multiple document comparison in the next section.
@@ 3.4 Single Document Summarization
4 Related Works
4.1 Surpass the Context Window Limitation
The quadratic time complexity $O(N^{2})$ of self-attention in transformer [1] poses limitations on training larger context window [24, 3]. A wide range of approaches has been explored to extend the context window of these models. One approach, exemplified by Longformer and its subsequent iterations [7], aims to handle much longer contexts by modifying the self-attention mechanism. Another strategy involves training models on shorter sequences and then modifying the Transformer’s structure to accommodate longer sequences during inference [25]. Lastly, some researchers are exploring methods such as chunking, leveraging discourse structures, and extracting key content to effectively reduce the context window required by LLMs, thereby facilitating the summarization of extensive documents [26]. Our paper will focus on the extraction method.
4.2 RAG & Long Document Summary
To give LLM the capability of utilizing unknown data, RAG is developed to retrieve relevant materials and utilize them for downstream tasks [4]. Several downstream approaches enhance RAG from various perspectives including graphRAG [27], correctiveRAG [28], hierarchical RAG [29], modular RAG [30] and more [31].
Long document summary includes extractive based method and abstractive based method. Initially, extractive methods, such as unsupervised Tf-Idf [32] and supervised BERT [2], were widely adopted until the advent of more potent decoder-based models. Subsequently, abstractive approaches gained prominence, exemplified by the end-to-end method [33] and hybrid extractive-abstractive strategies[34, 35, 36, 37].
Recently, Hypothetical Document Embedding (HyDE) applied LLM reranking to enhance diversity in Q&A tasks, claiming their method outperforms MMR, which provides nearly no benefit on downstream tasks [38, 39]. However, utilizing LLM for reranking can be expensive, slow and costly. In addition, these studies did not discuss the recall of relevant documents prior to LLM generation and explore various hyperparameters within MMR. This motivates us to explore more on MMR and FPS to enhance diversity, thereby improving the performance of LLMs on downstream tasks.
4.3 Representative Sentences Extraction
In the realm of extraction-based summarization, the criterion for choosing key sentences is crucial. Tf-Idf selected sentences based on the frequency of keyword [32]. PEGASUS selected sentences that utilizes structural information and ROUGE scores [34]. BERT selected representative sentences by distance in embedding space [2]. SentenceBERT [18], DeBERTa [40], E5 [19] and DPR [41] have been trained on larger corpora using more powerful computational resources. What is more, these methods only focus on selecting sentences that are related with core idea of the document, and it will cause problem for lacking the diversity of selected sentences. This research primarily focuses on leveraging smaller encoder-only models while incorporating diversity as a key consideration.
@@ 6 Limitation
There are several limitation on this works. To begin with, we only work on English dataset, while multilingual datasets should be tested to prove the importance of diversity on other language. In addition, this work focuses on research dataset while more work is supposed to be conducted on industrial datasets. Next, for extremely large dataset, more engineering work on parallelization like tree structures should be conducted to reduce latency. Finally, further research is needed to enhance LLMs’ ability to effectively utilize retrieved documents—such as through post-training methods that encourage the model to place greater emphasis on the retrieved information.
Appendix A Prompts for Q&A and Summarization
@@
Appendix B Impacts of Hyperparameters in Diversity for Retrieval
The impact of different hyperparameters: $\alpha$, $w$, $c_{r}$ on the recall of the Natural Question dataset of single document Q&A is shown in Figure. 6. The first and second subfigures illustrate the recall ratios of answers contained in the selected documents for SB+MMR and SB+FPS. When the weight parameter $w=1$, they are equivalent to SB. From the results, we can conclude that both SB+MMR and SB+FPS outperform SB. The last two subfigures display the latency of SB+MMR and SB+FPS. SB+FPS shows slightly worse performances than SB+MMR, and the latency of SB+MMR is significantly lower, especially when the context window is very long. Considering these two aspects, SB+MMR is more suitable for practical use compared to SB+FPS.
Figure 6: The impact of different hyperparameters: $\alpha$, $w$, $c_{r}$ on the recall of the Natural Question dataset of single document Q&A.